# Ensuring Safety in an Uncertain Environment: Constrained MDPs via Stochastic Thresholds

### v2

@@ Abstract
This paper studies constrained Markov decision processes (CMDPs) with constraints against stochastic thresholds, aiming at the safety of reinforcement learning in unknown and uncertain environments. We leverage a Growing-Window estimator sampling from interactions with the uncertain and dynamic environment to estimate the thresholds, based on which we design Stochastic Pessimistic-Optimistic Thresholding (SPOT), a novel model-based primal-dual algorithm for multiple constraints against stochastic thresholds. SPOT enables reinforcement learning under both pessimistic and optimistic threshold settings. We prove that our algorithm achieves sublinear regret and constraint violation; i.e., a reward regret of $\tilde{\mathcal{O}}(\sqrt{T})$ while allowing an $\tilde{\mathcal{O}}(\sqrt{T})$ constraint violation over $T$ episodes. The theoretical guarantees show that our algorithm achieves performance comparable to that of an approach relying on fixed and clear thresholds. To the best of our knowledge, SPOT is the first reinforcement learning algorithm that realises theoretical guaranteed performance in an uncertain environment where even thresholds are unknown.
@@ 1 Introduction
Reinforcement learning has enabled major advances in AI, such as AlphaGo silver2016mastering , autonomous vehicle control kendall2019learning , and LLM alignment ouyang2022training , in which an agent is trained to maximise its expected cumulative reward from the environment Sutton1998 . Markov decision processes (MDPs) are widely adopted in reinforcement learning, where transitions and rewards depend only on the current state rather than on the entire history bellman1957markovian . In many applications, ensuring safety is of profound importance – for instance, autonomous vehicles are expected to avoid collisions and adhere to traffic regulations wen2020safe . This motivates constrained Markov decision processes (CMDPs), introducing safety constraints against thresholds into optimisation of the cumulative reward function altman1999constrained .
Existing research has shown that linear-programming and primal-dual methods are able to learn a (near-)optimal policy for selecting actions from given states zheng2020constrained ; efroni2020exploration ; wei2018online ; qiu2020upper . However, these methods rely on a prevailing assumption that thresholds can be pre-determined, which is not true in real applications where environmental knowledge is highly limited, and thresholds are unknown and uncertain. Consider the example of autonomous vehicles – they can enter an unseen area where safety thresholds are unknown. Our vision is that an agent could have to infer thresholds from interactions with the environment. Hence, an open problem arises:
@@
Can we design a CMDP algorithm that achieves decent guarantees on its efficiency and safety in an environment with unknown safety thresholds?
@@
To address this problem, this paper proposes stochastic pessimistic-optimistic thresholding (SPOT), a new algorithm with theoretical guarantees on both safety and performance. Our primary objective is a strategy for CMDPs with finite horizon, similar to an array of existing works bai2020provably ; bura2022dope ; efroni2020exploration ; ghosh2024towards ; muller2024truly ; stradi2024optimal ; zheng2020constrained , but we consider the situation that the agent does not know the thresholds. The agent’s efficiency and safety are measured by two regrets defined with respect to (w.r.t.) the reward function and the constraint violations, respectively. We make contributions as follows:
@@
  • Mathematically grounding unknown and uncertain thresholds in CMDPs. This paper first realises and defines a safe reinforcement learning problem in which knowledge of the environment is highly limited that even the thresholds are unknown, which is termed as CMDP-ST (CMDP with stochastic thresholds).
@@
  • The first approach to safe reinforcement learning agents in such uncertain environment. SPOT leverages a Growing-Window estimator sampling during the agent’s interactions with the environment to estimate the thresholds. Based on this estimator, SPOT employs the primal-dual scheme to integrate the estimated thresholds into the learning process of the agent. SPOT is compatible with two variants that (1) employs pessimistic thresholds to strictly enforce safety, and (2) applies optimistic thresholds to encourage exploration and boost performance. By systematically navigating the latent ‘margin’ between safety and performance without relying on any prior threshold information, our approach offers a dynamic balance between exploitation and exploration.
  • Theoretical guarantees on threshold estimator and the induced optimisation problem. We first prove that our threshold estimator is asymptotically consistent with the hidden thresholds with a guaranteed convergence speed. Based on this estimator, we can then properly define a constrained optimisation problem for CMDP-ST. We also prove the feasibility and strong duality of this constrained optimisation problem, which respectively suggest that (1) at least one policy exists which meets all constraints, and (2) the optimal value of the original problem is maintained in its dual problem (Section 3.3).
@@
  • Theoretical guarantees on the efficiency and safety of SPOT. We prove that SPOT achieves $\tilde{\mathcal{O}}(\sqrt{T})$ regrets, for both cumulative reward and constraint violations over $T$ episodes without a pre-defined threshold, which are comparable to its counterpart with pre-defined thresholds (Section 4). Our analysis also suggests a trade-off between efficiency and safety: a constraint that enforces stringent safety can effectively mitigate violations but could incur higher reward regret, and vice versa.
@@
Related Works. Research on MDPs for online learning has seen a long history, which primarily focused on achieving a sublinear regret in stochastic or adversarial environments orabona2019modern ; auer2008near ; even2009online ; neu2010online ; rosenberg2019online ; rosenberg2019online2 ; jin2020learning . Building on these, researchers have integrated safety and resource constraints into MDPs, yielding CMDP methods that achieve strong sublinear regret bounds under assumptions of ‘full-information feedback’, where the complete loss function is observable wei2018online ; qiu2020upper . Subsequent works on bandit feedback, where the agent only observes the outcomes of its chosen actions, extended these guarantees to both known- and unknown-transition models, achieving sublinear bounds on both weak reward regret and constraint violation zheng2020constrained ; efroni2020exploration ; bai2020provably ; bura2022dope . Recent works have pushed the frontier further, obtaining sublinear strong regret guarantees stradi2024optimal ; muller2024truly ; ghosh2024towards . Efforts have also been seen in extending the results to more complex non-stationary environments (i.e., under assumptions that rewards and constraints have bounded variations), maintaining sublinear regrets even as the environment evolves ding2023provably ; wei2023provably . Although partial feedback models have been explored beyond simple bandit scenarios bacchiocchi2024markov ; lin2024information , these methods generally yield weaker theoretical guarantees. In addition, adversarial scenarios remain particularly challenging, with most existing results achieving only weak regret bounds stradi2024online ; levy2023efficient . As discussed, the entire literature assumes that safety thresholds can be clearly pre-defined. To our best knowledge, this work is the first to move beyond this limitation, studying safe reinforcement learning problem in an uncertain environment where thresholds are unknown.
@@ 2 Preliminaries
An MDP characterises the sequential decision-making process of an agent, positioned in a state $s$ from a finite state space $\mathcal{S}$ (with cardinality $|\mathcal{S}|=S$), which could select an action $a$ from a finite action space $\mathcal{A}$ (with cardinality $|\mathcal{A}|=A$) and then transition to another state $s^{\prime}\in\mathcal{S}$ according to a transition function $p:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$. The course of each episode starts at an initial state $s_{1}\in\mathcal{S}$ and continues for $H\in\mathbb{N}$ steps, which is named as horizon. At each step $h\in[H]$, the agent receives a non-stationary reward $\tilde{r}_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ drawn from a distribution $\mathcal{R}$ with expectation reward $r_{h}(s,a)$, which means its underlying distribution may vary over time. The agent’s trajectory is determined by a policy $\pi$ consisting of decision rules $(\pi_{1},\ldots,\pi_{H})\in\Pi$, where each $\pi_{h}(\cdot|s)\in\Delta(\mathcal{A})$ is a probability distribution over the action space $\mathcal{A}$. Here, $\Delta(\mathcal{A})$ denotes the probability simplex over the action space $\mathcal{A}$. A Constrained MDP (CMDP) extends an MDP by incorporating $m$ constraints, each defined by a function $g_{i,h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ associated with a threshold $\alpha_{i}\in[0,H]$. These thresholds collectively form a vector $\alpha\in[0,H]^{m}$. In the non-stationary scenario, the immediate constraint $\tilde{g}_{i,h}(s,a)\in[0,1]$ is drawn from a distribution $\mathcal{G}_{i}$ with expectation $\mathbb{E}_{\mathcal{G}_{i}}[\tilde{g}_{i,h}(s,a)]=g_{i,h}(s,a)$ for every $i$-th constraint. Overall, a CMDP can be represented by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},H,p,r,g,\alpha)$.
@@
capture the expected sum of $v$ from step $h$ onward. For brevity, let $V_{v}^{\pi}$ denote $V_{v,1}^{\pi}(s_{1})$, the expected return starting from the initial state $s_{1}$. The primary objective is to find a policy that maximises the expected cumulative reward while satisfying all the constraints; formally,
@@
The reward regret measures the cumulative gap between the rewards obtained by the selected policies and the optimal policy $\pi^{\star}\in\Pi$, and the constraint violation measures how much the constraints are violated during the learning process. The objective of the agent is to select policies $\{\pi_{t}\}_{t=1}^{T}$ such that both $\mathcal{R}_{T}(r)$ and $\mathcal{R}_{T}(g)$ grow sublinearly in the number of episodes $T$.
@@ 3.1 Unknown and Stochastic Environment in Terms of Thresholds
Unknown thresholds are common in real-world scenarios. A prime example is that autonomous vehicles may encounter unseen roads, or even travel to foreign countries with significantly different traffic laws, making their safety thresholds unknown. Even on the same road, the safety thresholds for collision avoidance can be unclear due to extreme weather, poor visibility, or unexpected traffic incidents lindner2024learning . A similar situation arises in healthcare, such as dose allocation in clinical trials. For a new drug, safety thresholds are certainly unknown shen2020learning .
@@
Motivated by the examples above, we propose training a CMDP in which thresholds are unknown, and must therefore be inferred through real-time sampling. Specially, for each constraint $i\in[m]$, the threshold $\tilde{\alpha}_{i,h}(s,a)\in[0,H]$ for any state-action pair $(s,a)$ is drawn from a distribution $\mathcal{L}_{i}$ with support of $[0,H]^{\mathcal{S}\times\mathcal{A}}$, whose expectation is $\mathbb{E}_{\mathcal{L}_{i}}[\tilde{\alpha}_{i,h}(s,a)]=\alpha_{i,h}(s,a)$.
Algorithm 1 depicts how an agent interacts with the environment, when rewards, constraints and thresholds are all stochastic. At each episode $t\in[T]$, the agent chooses a policy $\pi_{t}$, while independently sampling reward $r^{t}\sim\mathcal{R}$, constraint $g_{i}^{t}\sim\mathcal{G}_{i}$ and threshold $\alpha_{i}^{t}\sim\mathcal{L}_{i}$ for the $i$-th constraint. It is important to note that the agent has no prior knowledge of the reward, constraint, and threshold distributions, nor the transition function $p$.
@@
0:  Policy $\pi_{t}\in\Pi$, $\tilde{r}^{t}\in[0,1]^{\mathcal{S}\times\mathcal{A}}$, $\tilde{g}_{i}^{t}\in[0,1]^{\mathcal{S}\times\mathcal{A}}$, $\tilde{\alpha}_{i}^{t}\in[0,H]^{\mathcal{S}\times\mathcal{A}}$ for $i\in[m]$
@@
3:     Agent takes action $a\sim\pi_{t}(\cdot\mid s)$
4:     Agent observes reward $\tilde{r}_{h}^{t}(s,a)$, constraint $\tilde{g}_{i,h}^{t}(s,a)$ and threshold $\tilde{\alpha}_{i,h}^{t}(s,a)$ for $i\in[m]$
5:     Environment evolves to $s\sim p(\cdot\mid s,a)$
Dynamical nature of thresholds
Also motivated by the examples, we note that the environment is not static – rewards, constraints, and thresholds change over time. To take these into account, we define that rewards, constraints and thresholds are drawn from dynamical distributions $\mathcal{R}_{t}$, $\{\mathcal{G}_{i,t}\}_{i=1}^{m}$, $\{\mathcal{L}_{i,t}\}_{i=1}^{m}$ over $T$ episodes, instead of fixed distributions. Consequently, their expectations $r_{h}^{t}(s,a)$, $g_{i,h}^{t}(s,a)$, $\alpha_{i,h}^{t}(s,a)$ also evolve over time.
Occupancy measure
Given a policy $\pi$ and transition function $p$, the occupancy measure $q_{h}^{\pi}$ is defined as the probability of the agent visiting state $s$ and taking action $a$ at step $h$, that is, $q_{h}^{\pi}(s,a;p)=\mathbb{P}\left\{s_{h}=s,a_{h}=a|p,\pi\right\}$ altman1999constrained ; Puterman1994 . The entire collection of these probabilities across all steps, states and actions is denoted by $q^{\pi}(p):=\{q_{h}^{\pi}(s,a;p)\}_{H\times\mathcal{S}\times\mathcal{A}}$.
Probability measure of thresholds and feasible set
We introduce a probability space $(\Omega,\mathcal{F},\mathbb{P})$ to capture the randomness of thresholds, where each outcome $\omega\in\Omega$ represents a specific realisation of these values, referred to as a threshold scenario. For every constraint $i\in[m]$, the threshold $\alpha_{i}(\omega_{i})$ is drawn from an independent probability space $(\Omega_{i},\mathcal{F}_{i},\mathbb{P}_{i})$. Consequently, the overall threshold vector is $\alpha=(\alpha_{1}(\omega_{1}),\ldots,\alpha_{m}(\omega_{m}))$, where the threshold scenario $\omega=(\omega_{1},\ldots,\omega_{m})$. Under the occupancy measure formulation, the optimisation problem then becomes
$\max_{q\in\Delta{(\mathcal{M})}}\;\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\mathcal{L}(q,\lambda,\omega):=r^{T}q^{\pi}(p)+\lambda^{T}(G^{T}q^{\pi}(p)-\alpha(\omega)).$
where $G$ denotes the constraint matrix induced by the values $g_{i,h}(s,a)$ for all $i\in[m]$, $h\in[H]$, and $(s,a)\in\mathcal{S}\times\mathcal{A}$. Moreover, the corresponding stochastic feasible set is
$\mathcal{F}_{q}(\omega):=\{q\in\Delta{(\mathcal{M})}|\sum_{h,s,a}q_{h}^{\pi}(s,a;p)g_{i,h}(s,a)\geq\alpha_{i}(\omega_{i}),\forall i\in[m]\},$
which forms a random polytope within the space of occupancy measures $\Delta(\mathcal{M})$. Finally, the complete probability space for all constraints is defined as $(\Omega_{1}\times\cdots\times\Omega_{m},\mathcal{F}_{1}\otimes\cdots\otimes\mathcal{F}_{m},\mathbb{P}_{1}\otimes\cdots\otimes\mathbb{P}_{m})$.

### v3

@@ Abstract
This paper studies constrained Markov decision processes (CMDPs) with constraints against stochastic thresholds, aiming at safety of reinforcement learning in unknown and uncertain environments. We leverage a Growing-Window estimator sampling from interactions with the uncertain environment to estimate the thresholds, based on which we design Stochastic Pessimistic-Optimistic Thresholding (SPOT), a novel model-based primal-dual algorithm for multiple constraints against stochastic thresholds. SPOT enables reinforcement learning under both pessimistic and optimistic threshold settings. We prove that our algorithm achieves sublinear regret and constraint violation, i.e., a reward regret of $\tilde{\mathcal{O}}(\sqrt{T})$ while allowing an $\tilde{\mathcal{O}}(\sqrt{T})$ constraint violation over $T$ episodes. The theoretical guarantees show that our algorithm achieves performance comparable to that of an approach relying on fixed and clear thresholds. To the best of our knowledge, SPOT is the first reinforcement learning algorithm that realises theoretical guaranteed performance in an uncertain environment where even thresholds are unknown.
@@ 1 Introduction
Reinforcement learning (RL) has enabled major advances in AI, such as AlphaGo silver2016mastering, autonomous vehicle control kendall2019learning, and LLM alignment ouyang2022training, where an agent is trained to maximise its expected cumulative reward from the environment Sutton1998. Markov decision processes (MDPs) are widely adopted in reinforcement learning, where transitions and rewards depend only on the current state rather than on the entire history bellman1957markovian. In many applications, ensuring safety is of profound importance – for instance, autonomous vehicles are expected to avoid collisions and adhere to traffic regulations wen2020safe. This motivates constrained Markov decision processes (CMDPs), introducing safety constraints against thresholds into optimisation of the cumulative reward function altman1999constrained.
A substantial literature studies unknown constraints in safe RL. Classical CMDP methods, built on linear programming and primal dual updates, achieve (near-)optimal learning with regret guarantees under model uncertainty zheng2020constrained; efroni2020exploration; wei2018online; qiu2020upper. To reduce reliance on fully specified constraints, inverse constrained RL (ICRL) infers latent constraints or safety surrogates from demonstrations and can generalize across thresholds malik2021inverse; liu2024comprehensive; in parallel, constraint-conditioned approaches explicitly condition policies on a threshold input to enable cross-threshold generalization yao2023constraint. Safety-first exploration takes a different route: it does not estimate the threshold value themselves but enforces safety by learning certified safe sets wachi2020safe or by using uncertainty to remain below a specified safety budget wachi2023safe. Complementary lines manage budgets and risks through state augmentation sootla2022saute, chance constraints with randomized policies shen2024flipping and offline adaptation schemes that switch among pretrained policies to satisfy deployment-time cost thresholds chemingui2025constraint.
Despite this progress, most safe-RL methods still hinge on exogenous and fixed safety thresholds, rely on high-quality demonstrations, or substitute true limits with conservative surrogates wachi2024survey. These choices fail when admissible limits are numerically unknown and fluctuate with noise or operating conditions yao2023constraint. Our vision is to move beyond ‘adaptation around a given threshold’ and instead learn the threshold itself as a latent, data-driven quantity from interaction and update it online with calibrated uncertainty. This motivates the following question:
@@
Can we design an online CMDP algorithm that achieves decent theoretical guarantees in an environment with unknown safety thresholds?
@@
To address this problem, this paper proposes stochastic pessimistic-optimistic thresholding (SPOT), an online algorithm with theoretical guarantees on both safety and performance when thresholds evolve over time. We consider CMDPs with finite horizon, similar to an array of existing works bai2020provably; bura2022dope; efroni2020exploration; ghosh2024towards; muller2024truly; stradi2024optimal; zheng2020constrained. We make contributions as follows:
@@
  • Mathematically grounding stochastic thresholds in CMDPs. This paper first realises and defines a safe reinforcement learning problem in which knowledge of the environment is highly limited that even the thresholds are unknown, which is termed as CMDP-ST (CMDP with stochastic thresholds). We formalise this uncertainty by modelling the threshold as a stochastic process: instead of observing the true episodic safety thresholds ($\alpha_{i}$), the agent only receives noisy and per-step signals ($\tilde{\alpha}_{i,h}$), where the latent threshold is defined as their expected sum. This grounds the problem of learning to satisfy constraints online without any prior knowledge of the threshold’s underlying distribution.
@@
  • The first approach to safe reinforcement learning agents in such uncertain environment. SPOT leverages a Growing-Window estimator that, at each step and episode, uses the value of the most visited state-action pair to characterize the global threshold estimate. Based on this estimator, SPOT employs a primal-dual scheme to integrate the thresholds into the learnt agent. SPOT is compatible to two variants that (1) employs pessimistic thresholds to strictly enforce safety, and (2) applies optimistic thresholds to encourage exploration. By systematically navigating the latent ‘margin’ between safety and performance without relying on any prior threshold information, our approach offers a dynamic balance between exploitation and exploration.
@@
  • Theoretical guarantees on the efficiency and safety of SPOT. We first prove that our threshold estimator is asymptotically consistent with the hidden thresholds with a guaranteed convergence speed (Section 3.2). We then establish the feasibility and strong duality of this constrained optimisation problem (Section 3.3). We prove that SPOT achieves $\tilde{\mathcal{O}}(\sqrt{T})$ regrets, for both cumulative reward and constraint violations over $T$ episodes, which are comparable to its counterpart with pre-defined thresholds (Section 4). Our analysis also suggests a trade-off between efficiency and safety: a constraint that enforces stringent safety can effectively mitigate violations but incur higher reward regret, and vice versa.
@@
Related Works. Research on MDPs for online learning has seen a long history, which primarily focused on achieving a sublinear regret in stochastic or adversarial environments orabona2019modern; auer2008near; even2009online; neu2010online; rosenberg2019online; rosenberg2019online2; jin2020learning. Building on these, researchers have integrated safety and resource constraints into MDPs, yielding CMDP methods that achieve strong sublinear regret bounds under assumptions of ‘full-information feedback’, where the complete loss function is observable wei2018online; qiu2020upper. Subsequent works on bandit feedback, where the agent only observes the outcomes of its chosen actions, extended these guarantees to both known- and unknown-transition models, achieving sublinear bounds on both weak reward regret and constraint violation zheng2020constrained; efroni2020exploration; bai2020provably; bura2022dope. Recent works have pushed the frontier further, obtaining sublinear strong regret guarantees stradi2024optimal; muller2024truly; ghosh2024towards. Efforts have also been seen in extending the results to more complex non-stationary environments (i.e., under assumptions that rewards and constraints have bounded variations), maintaining sublinear regrets even as the environment evolves ding2023provably; wei2023provably. Although partial feedback models have been explored beyond simple bandit scenarios bacchiocchi2024markov; lin2024information, these methods generally yield weaker theoretical guarantees. In addition, adversarial scenarios remain particularly challenging, with most existing results achieving only weak regret bounds stradi2024online; levy2023efficient. As discussed, most studies assume that safety thresholds can be clearly pre-defined. To our best knowledge, this paper presents the first CMDP algorithm to move beyond this limitation, learning safety thresholds online from interaction while achieving sublinear regret bounds.
@@ 2 Preliminaries
An MDP characterises the sequential decision-making process of an agent, positioned in a state $s$ from a finite state $\mathcal{S}$ (with $|\mathcal{S}|=S$), which could select an action $a$ from a finite action space $\mathcal{A}$ (with $|\mathcal{A}|=A$) and then transition to another state $s^{\prime}\in\mathcal{S}$ according to a transition function $p_{h}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to[0,1]$. The course of each episode starts at an initial state $s_{1}\in\mathcal{S}$ and continues for $H\in\mathbb{N}$ steps, which is named as horizon. At each step $h\in[H]$, the agent receives a reward $\tilde{r}_{h}:\mathcal{S}\times\mathcal{A}\to[0,1]$ drawn from a distribution $\mathcal{R}_{h}$ with expectation reward $r_{h}(s,a)$, which means its underlying distribution may vary over time. The agent’s trajectory is determined by a policy $\pi$ consisting of decision rules $(\pi_{1},\ldots,\pi_{H})\in\Pi$, where each $\pi_{h}(\cdot|s)\in\Delta(\mathcal{A})$ is a probability distribution over the action space $\mathcal{A}$. Here, $\Delta(\mathcal{A})$ denotes the probability simplex over the action space $\mathcal{A}$. A Constrained MDP (CMDP) extends an MDP by incorporating $m$ constraints, each defined by a function $g_{i,h}:\mathcal{S}\times\mathcal{A}\to[0,1]^{m}$ associated with a threshold. These thresholds collectively form a vector $\alpha\in[0,H]^{m}$. In the stationary scenario, the immediate constraint $\tilde{g}_{i,h}(s,a)\in[0,1]$ is drawn from a distribution $\mathcal{G}_{i,h}$ with expectation $\mathbb{E}_{\mathcal{G}_{i}}[\tilde{g}_{i,h}(s,a)]=g_{i,h}(s,a)$ for every constraint $i\in[m]$. Overall, a CMDP can be represented by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},H,p,r,g,\alpha)$.
@@
capture the expected sum of $v$ from step $h$ onward. For brevity, let $V_{v}^{\pi}$ denote $V_{v,1}^{\pi}(s_{1})$, the expected return starting at the initial state. The primary objective is to find a policy that maximises the expected cumulative reward while satisfying all the constraints; formally,
@@
The reward regret measures the cumulative gap between the rewards obtained by the selected policies and the optimal policy $\pi^{\star}\in\Pi$, and the constraint violation measures how much the constraints are violated during the learning process.
@@ 3.1 Unknown and Stochastic Environment in Terms of Thresholds
In many real-world applications, safety or performance thresholds are not pre-defined but must be inferred from a stream of noisy observations. A representative example is personalized medicine: a treatment plan for a patient might span several weeks and total cumulative dose of a drug must stay below a toxicity level for the entire period. At each step, for instance on a daily basis, a doctor receives noisy signals from the patient’s biomarkers. This feedback provides a per-step hint about the safe dosage for that day liu2024learning. Therefore, the agent’s core challenge is to aggregate these per-step, unreliable signals to form a coherent strategy for the overall episodic constraint.
@@
Motivated by this example, we propose training a CMDP in which the safety thresholds are realized through a stochastic process. We assume a stationary environment in which, for each constraint $i\in[m]$ and step $h\in[H]$, there exists an unknown threshold $\alpha_{i,h}\in[0,1]$. When the agent visits any state-action pair $(s,a)$, it receives a bounded noisy observation $\tilde{\alpha}_{i,h}(s,a)\in[0,1]$ which is drawn from a distribution $\mathcal{L}_{i,h}$ with expectation $\mathbb{E}_{\mathcal{L}_{i,h}}[\tilde{\alpha}_{i,h}(s,a)]=\alpha_{i,h}$. The episodic threshold for constraint $i$ is $\alpha_{i}:=\sum_{h=1}^{H}\alpha_{i,h}$. Similarly, for each step $h$ and constraint $i$, the reward $\tilde{r}_{h}(s,a)$ and cost $\tilde{g}_{i,h}(s,a)$ are drawn from unknown families of distributions $\mathcal{R}_{h}$ and $\mathcal{G}_{i,h}$, respectively.
The fundamental difficulty is that the agent has no prior knowledge of the distributions $\{\mathcal{L}_{i,h}\}_{i,h}$, $\{\mathcal{R}_{h}\}_{h}$, or $\{\mathcal{G}_{i,h}\}_{i,h}$. Consequently, the true parameters which involve the expected reward vector $r$, constraints $g_{i}$ and thresholds $\alpha_{i}$ are all unknown. This information must be acquired online by sampling from these distributions. Algorithm 1 depicts this interaction, where at each episode $t$, the agent executes a policy $\pi_{t}$ and observes stochastic realisations (i.e., the samples $\tilde{r}_{h}$, $\tilde{g}_{i,h}$ and $\tilde{\alpha_{i,h}}$, which are then used to learn the unknown parameters.
@@
0: Policy $\pi_{t}\in\Pi$
@@
3:  Agent takes action $a_{h}\sim\pi_{t}(\cdot\mid s_{h})$
4:  Agent observes reward $\tilde{r}_{h}^{t}(s_{h},a_{h})$, constraint cost $\tilde{g}_{i,h}^{t}(s_{h},a_{h})$, and threshold $\tilde{\alpha}_{i,h}^{t}(s_{h},a_{h})$ for all $i\in[m]$
5:  Environment evolves to $s_{h+1}\sim p(\cdot\mid s_{h},a_{h})$
Objective Function under Stochastic Thresholds The learning objective is to find a policy $\pi^{\star}$ that maximizes the expected cumulative reward while respecting the (unknown) safety constraints:
$\max_{\pi\in\Pi}V_{r}^{\pi}\quad\text{s.t.}\quad V_{g_{i}}^{\pi}\geq\alpha_{i}=\sum_{h=1}^{H}\alpha_{i,h},\quad\forall i\in[m],$
where $V_{r}^{\pi}:=\mathbb{E}_{\pi,p}\left[\sum_{h=1}^{H}r_{h}\right]$ and $V_{g_{i}}^{\pi}:=\mathbb{E}_{\pi,p}\left[\sum_{h=1}^{H}g_{i,h}\right]$ are the reward and constraint expectations under the trajectory induced by $(\pi,p)$, and $\alpha_{i}=\sum_{h=1}^{H}\alpha_{i,h}$ is the threshold expectation with $\alpha_{i,h}=\mathbb{E}_{\mathcal{L}_{i,h}}[\tilde{\alpha}_{i,h}]$.
Given a policy $\pi$ and transition function $p$, the occupancy measure $q_{h}^{\pi}$ is defined as the probability of the agent visiting state $s$ and taking action $a$ at step $h$, that is, $q_{h}^{\pi}(s,a;p)=\mathbb{P}\left\{s_{h}=s,a_{h}=a|p,\pi\right\}$ Puterman1994; altman1999constrained. The entire collection of these probabilities is denoted by $q:=[q_{h}^{\pi}(s,a;p)]_{H\times\mathcal{S}\times\mathcal{A}}$. Define the reward vector $r$ with entries $r_{h}(s,a)$ and the constraint matrix $G$ with rows $(g_{i,h}(s,a))_{h,s,a}$. Then, $V_{r}^{\pi}=r^{\top}q$ and $(V_{g_{i}}^{\pi})_{i\in[m]}=G^{\top}q$. Under the occupancy measure formulation, the optimisation problem is a saddle-point problem defined over the (unknown) expected reward vector $r$, constraint matrix $G$ and threshold vector $\alpha$:
$\max_{q\in\Delta{(\mathcal{M})}}\;\min_{\lambda\in\mathbb{R}^{m}_{\geq 0}}\mathcal{L}(q,\lambda):=r^{\top}q+\lambda^{\top}(G^{\top}q-\alpha).$
The corresponding target feasible set is $\mathcal{F}_{q}:=\{q\in\Delta{(\mathcal{M})}|\sum_{h,s,a}q_{h}^{\pi}(s,a;p)g_{i,h}(s,a)\geq\alpha_{i},\forall i\in[m]\}$, which forms a convex polytope within the space of occupancy measures $\Delta(\mathcal{M})$.